A)A pair of oppositely charged parallel plates is separated by 5.52 mm. A potential difference of 610 V exists between the plates. What is the strength of the electric field between the plates? The fundamental charge is 1.602 × 10−19 . Answer in units of V/m.

B)What is the magnitude of the force on an electron between the plates? Answer in units of N.

C)How much work must be done on the electron to move it to the negative plate if it is initially positioned 2.88 mm from the positive plate? Answer in units of J.
Answer #1

A)

Electric field, E = V/d = 610/0.00552 = 1.11 x 10^5 N/C

B)

Force, F = qE = (1.6 x 10^-19)(1.11 x 10^5) = 1.77 x 10^-14 N

C)

Work done, W = Fs = 1.77 x 10^-14 x (0.00552 - 0.00288) = 4.67 x 10^-17 J

Comment in case any doubt please rate my answer ....